# A generalization of the Choi isomorphism with application to open quantum systems

### v1

@@ I Introduction
It has proved useful to use a general concept of “state changes" in quantum mechanics, which includes time evolutions of closed and open systems, as well as the changes due to measurements which generalize von Neumann’s projection postulate. Since (pure or mixed) quantum states are represented by statistical operators (density matrices), it is natural to describe such state changes by linear mappings that preserve the trace and map positive operators to positive ones.
@@
In the years that followed, the term “operation" was introduced for certain state changes in the work of Haag and Kastler HK64 and further elaborated by Davies and Lewis DL70 . The first steps towards a restriction of the general concept of an “operation" were taken by Hellwig and Kraus in HK69 and HK70 , who developed their approach under the influence of Ludwig’s axiomatic foundation of quantum mechanics L70 . They derived what is known today as the “operator-sum representation" or “Kraus operator reresentation" of an operation, see, e.g., NC00 , and which has been recognized as being equivalent to complete positivity in K71 . These results are usually quoted from the “Lecture notes in physics" volume K83 .
After this breakthrough, it is not surprising that, as already alluded to in SMR61 , completely positive mappings are also used in the time evolution of open quantum systems. In 1976 two papers appeared simultaneously investigating semigroups of completely positive time evolutions, GKS76 and L76 , which led to the so-called Lindblad equation. This term has become commonplace, though perhaps not entirely fair. For our purposes, eq. (2.4) in GKS76 is particularly interesting, since it establishes a 1:1 relationship between linear maps ${\mathcal{E}}$ of state spaces and Hermitean matrices, called GKS matrices $g$ in this paper, independent of the semigroup property. Consequently, the approach of GKS76 was later used to derive a time-dependent Lindblad equation which holds for general open quantum systems, see B12 or subsection VI.2 of this paper. We will prove that ${\mathcal{E}}$ is completely positive iff $g$ is positively semi-definite, thus generalizing Choi’s criterion, since the Choi matrix can be seen as a special case of the GKS matrix.
@@
These calculations do not yet assume that the GKS equation follows from the reduced interaction of the system $S$ with an environment $E$. In the subsection VI.3, however, we make this assumption and try to calculate the GKS matrix as a function of the Hamiltonian for the total system $S+E$. Due to the inherent difficulty of these calculations, we restrict ourselves to a series expansion up to the $2$nd order w. r. t. time $t$. Note that in the dynamic semigroup approximation, the reduced time derivative of the statistical operator $\rho(t)$ would be constant and thus its $2$nd derivative would vanish. Our calculations therefore lead to results that go beyond the semigroup approximation. We will focus on the question of whether the GKS matrix is positively semi-definite in $2$nd order in $t$. On the one hand, this follows in general since the time evolution of $\rho(t)$ is completely positive. On the other hand, it is instructive to show this directly in order to check the consistency of the present approach. We conclude with a summary in Section VII.
@@ II Definitions and general results
where $\mbox{T}r$ denotes the trace and $X^{\ast}$ the adjoint of the operator $X$. Further let ${\sf B}({\mathcal{H}})$ denote the real subspace of $L({\mathcal{H}})$ consisting of Hermitean operators, ${\sf B}_{0}({\mathcal{H}})\subset{\sf B}({\mathcal{H}})$ the subset of operators with vanishing trace, ${\sf B}^{+}({\mathcal{H}})\subset{\sf B}({\mathcal{H}})$ the cone consisting of positively semi-definite operators, ${\sf B}^{+}_{1}({\mathcal{H}})\subset{\sf B}^{+}({\mathcal{H}})$ the convex set of statistical operators. and ${\sf U}({\mathcal{H}})$ the group of unitary operators acting on ${\mathcal{H}}$. There exists a canonical isomorphism $\imath:{\mathcal{A}}\otimes{\mathcal{A}}\rightarrow L({\mathcal{H}}\otimes{\mathcal{H}})$ given by
@@
$\langle{\mathcal{X}}\left|\right.{\mathcal{Y}}\rangle:={\mathcal{T}r}\left({\mathcal{X}}^{\ast}\,{\mathcal{Y}}\right),\;{\mathcal{X}},{\mathcal{Y}}\in L({\mathcal{A}})\;.$ (4)
@@
We will formulate the definitions concerning “complete positivity" only for the special case of a linear map $\phi:{\mathcal{A}}\rightarrow{\mathcal{A}}$ considered in this article, following NC00 and C06 . Recall that $L\left({\mathbb{C}}^{M}\right)$ denotes the algebra of complex $M\times M$-matrices. $\phi$ can be extended to $\phi\otimes{\mathbbm{1}}_{N}:{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)\rightarrow{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)$ by $(\phi\otimes{\mathbbm{1}})(A\otimes M):=\phi(A)\otimes M$ and linear extension. Then $\phi$ is called “completely positive" iff $X\geq 0$ implies $(\phi\otimes{\mathbbm{1}}_{N})(X)\geq 0$ for all Hermitean operators $X\in{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)$. This condition is stronger than the property that $\phi$ maps positively semi-definite operators onto positively semi-definite ones. The usual example to illustrate this fact is $\phi(A)=A^{\top}$ (transposition). $\phi$ leaves the eigenvalues of any positively semi-definite operator invariant but is not completely positive: Set $N=2$ and ${\mathcal{A}}=L\left({\mathbb{C}}^{2}\right)$ and consider the matrix $X\in{\mathcal{A}}\otimes L\left({\mathbb{C}}^{2}\right)$ given by
@@
with eigenvalues $(2,0,0,0)$. $\phi\otimes{\mathbbm{1}}_{2}$ swaps the two anti-diagonal blocks of $X$
$(\phi\otimes{\mathbbm{1}}_{2}((X)=\left(\begin{array}[]{cc|cc}1&0&0&0\\ 0&0&1&0\\ \hline\cr 0&1&0&0\\ 0&0&0&1\\ \end{array}\right)\;,$ (6)
and thus yields a matrix with eigenvalues $(-1,1,1,1)$.
@@ IV.1 Choi isomorphism as a special case of the GKS isomorphism
It is shown in GKS76 that this family actually defines a basis of $L({\mathcal{A}})$ but we will not use this fact at the moment and postpone its proof. Let us first focus on the special case that an orthonormal basis $\left(|i\rangle\right)_{0\leq i<N}$ of ${\mathcal{H}}$ is chosen and consider the derived orthonormal basis $\left(|j\rangle\,\langle i|\right)_{0\leq i,j<N}$ of ${\mathcal{A}}$. We write $\alpha\equiv(i,j)$ as a double index and set $E_{\alpha}:=|j\rangle\,\langle i|$, hence $E_{\alpha}^{\ast}:=|i\rangle\,\langle j|$. Note the reversed positions of the indices. Then it can be shown that
@@
$\left\langle{\mathcal{E}}_{\alpha\beta}\left|\right.{\mathcal{E}}_{\gamma\delta}\right\rangle$ $=$ ${\mathcal{T}r}\left({\mathcal{E}}_{\alpha\beta}^{\ast}\,{\mathcal{E}}_{\gamma\delta}\right)$ (20)
@@
$=$ $\sum_{\epsilon}\mbox{Tr }\left(E_{\beta}\,E_{\epsilon}^{\ast}\,E_{\alpha}\,E_{\gamma}\,E_{\epsilon}\,E_{\delta}^{\ast}\right)$ (24)
@@
Then the matrix with entries $\widetilde{c}_{\alpha\beta}$ is exactly the Choi matrix of ${\mathcal{E}}$. More precisely, we have
@@ IV.2 Definition of the GKS isomorphisms
    $=$ $\sum_{\alpha^{\prime}\,\beta^{\prime}}\overline{U_{\alpha\alpha^{\prime}}}\,U_{\beta\beta^{\prime}}\,\delta_{\alpha^{\prime}\,\beta^{\prime}}=\sum_{\alpha^{\prime}\,\beta^{\prime}}\overline{U_{\alpha\alpha^{\prime}}}\,U_{\beta\alpha^{\prime}}=\sum_{\alpha^{\prime}\,\beta^{\prime}}U_{\beta\alpha^{\prime}}\,U^{\ast}_{\alpha^{\prime}\,\alpha}\;.$ (37)
@@
$G$ ::: $L({\mathcal{A}})\rightarrow L({\mathbb{C}}^{N\times N}),$ (44)
@@
The comparison of (42) and (28) shows that the Choi matrix isomorphism is a special case of the GKS-isomorphism. Already the Choi matrix isomorphism is not a single mathematical object but rather a family since it depends on a family of orthonormal bases in ${\mathcal{A}}$ that are of the form $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ w. r. t. the family of orthonormal bases in ${\mathcal{H}}$ denoted by $\left(|i\rangle\right)_{0\leq i<N}$. The GKS-isomorhism $G$ further extends this family by considering general orthonormal bases $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in ${\mathcal{A}}$. Of course, it still has to be shown that $G$ has analogous properties as the Choi matrix isomorphism.
@@ IV.3 Properties of the GKS isomorphisms
=:absent:=:= : $\sum_{\alpha^{\prime}\beta^{\prime}}\hat{g}^{\alpha^{\prime}\beta^{\prime}}\,G_{\alpha^{\prime}}\,A\,G_{\beta^{\prime}}^{\ast}\;.$ (48)
@@
Tr $\mbox{Tr }A$ $=$ $\mbox{Tr }{\mathcal{E}}(A)=\sum_{\alpha\beta}g^{\alpha\beta}\,\mbox{Tr }\left(F_{\alpha}\,A\,F_{\beta}^{\ast}\right)\quad\mbox{for all }A\in{\mathcal{A}}$ (49)
$\Leftrightarrow$ $\mbox{Tr }\left(\left(\sum_{\alpha\beta}g^{\alpha\beta}\,F_{\beta}^{\ast}\,F_{\alpha}\right)-{\mathbbm{1}}\right)A=0\quad\mbox{for all }A\in{\mathcal{A}}$ (50)
@@
=:absent:=:= : $\sum_{\gamma}g^{\gamma}\,A_{\gamma}\,X\,A_{\gamma}^{\ast}\;.$ (57)
@@ V Comparison with other isomorphisms
There exist various other generalizations of the Choi isomorphism and the question arises whether the GKS-isomorphism considered above is really different. We will check this by an application to the example of the transposition $\phi$ of $2\times 2$-matrices considered in Section II which maps positive matrices to positive ones but is not completely positive.
@@
in ${\mathcal{A}}$. Since $\phi$ only swaps $E_{1}$ and $E_{2}$ we can represent it by the matrix
$\widetilde{\phi}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{array}\right)\;.$ (60)
According to (17) the Choi matrix $c(\widetilde{\phi})$ of $\widetilde{\phi}$ is obtained by a certain permutation of matrix elements, see Figure 1. It follows that $c(\widetilde{\phi})=\widetilde{\phi}$ and hence $c(\widetilde{\phi})$ is not positively semi-definite in accordance with the fact that $\phi$ is not completely positive.
@@
in ${\mathcal{A}}$. After some calculations we obtain the following GKS-matrix $G(\phi)$
$G(\phi)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&1\\ \end{array}\right)\;,$ (62)
@@ V.1 Comparison with dePillis-Jamiołski isomorphism
As already noted in FC24 , the isomorphism due to dP67 and J72 differs from the Choi isomorphism, although this difference is sometimes blurred by the misleading denotation “Choi-Jamiołski isomorphism". It can be written in the form
@@
seemingly depending on an orthonormal basis $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in ${\mathcal{A}}$, but $J({\mathcal{E}})$ is be shown to be independent of this basis, see FC24 . Again we calculate the matrix of $J(\phi)$ where $\phi$ denotes the transposition of $2$-matrices and obtain
$J(\phi)=\left(\begin{array}[]{cccc}1&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&1\\ \end{array}\right)\;,$ (64)

### v2

@@ I Introduction
It has proved useful to employ a general concept of “state changes" in quantum mechanics, which includes time evolutions of closed and open systems, as well as the changes due to measurements which generalize von Neumann’s projection postulate. Since (pure or mixed) quantum states are represented by statistical operators (density matrices), it is natural to describe such state changes by linear mappings that preserve the trace and map positive operators to positive ones.
@@
In the years that followed, the term “operation" was introduced for certain state changes in the work of Haag and Kastler HK64 and further elaborated by Davies and Lewis DL70 . The first steps towards a restriction of the general concept of an “operation" were taken by Hellwig and Kraus in HK69 and HK70 , who developed their approach under the influence of Ludwig’s axiomatic foundation of quantum mechanics L70 . They derived what is known today as the “operator-sum representation" or “Kraus operator representation" of an operation, see, e.g., NC00 , and which has been recognized as being equivalent to complete positivity in K71 . These results are usually quoted from the “Lecture notes in physics" volume K83 .
After this breakthrough, it is not surprising that, as already alluded to in SMR61 , completely positive mappings are also used in the time evolution of open quantum systems. In 1976 two papers appeared simultaneously investigating semigroups of completely positive time evolutions, GKS76 and L76 , which led to the so-called Lindblad equation. This term has become commonplace, though perhaps not entirely fair. It is interesting for the history of the Lindblad equation that the first author of GKS76 , V. Gorini, became acquainted with the concept of completely positive mappings in discussions with K. Kraus at a conference in Marburg in 1973 HN74 . He told this to A. Kossakowski and informed E. Sudarshan, who invited to Austin both Kossakowski and Gorini for a few months in 1975-76, where the paper GKS76 was written VG24 . This part of the history of the Gorini-Kossakowski-Sudarshan-Lindblad equation is also touched in W25 .
For our purposes, eq. (2.4) in GKS76 is particularly interesting, since it establishes a 1:1 relationship between linear maps $\mathcal{E}$} of state spaces and Hermitean matrices, called GKS matrices $g$ in this paper, independent of the semigroup property. Consequently, the approach of GKS76 was later used to derive a time-dependent Lindblad equation which holds for general open quantum systems, see B12 or subsection VI.2 of this paper. We will prove that $\mathcal{E}$} is completely positive iff $g$ is positively semi-definite, thus generalizing Choi’s criterion, since the Choi matrix can be seen as a special case of the GKS matrix.
@@
These calculations do not yet assume that the GKS equation follows from the reduced interaction of the system $S$ with an environment $E$. In the subsection VI.3, however, we make this assumption and try to calculate the GKS matrix as a function of the Hamiltonian for the total system $S+E$. Due to the inherent difficulty of these calculations, we restrict ourselves to a series expansion up to the 22nd order w. r. t. time $t$. Note that in the dynamic semigroup approximation, the reduced time derivative of the statistical operator $\rho(t)$ would be constant and thus its 22nd derivative would vanish. Our calculations therefore lead to results that go beyond the semigroup approximation. We will focus on the question of whether the GKS matrix is positively semi-definite in 22nd order in $t$. On the one hand, this follows in general since the time evolution of $\rho(t)$ is completely positive. On the other hand, it is instructive to show this directly in order to check the consistency of the present approach. We conclude with a summary in Section VII.
@@ II Definitions and general results
where Tr denotes the trace and $X^{\ast}$ the adjoint of the operator $X$. Further let ${\sf B}({\mathcal{H}})$ denote the real subspace of $L({\mathcal{H}})$ consisting of Hermitean operators, ${\sf B}_{0}({\mathcal{H}})\subset{\sf B}({\mathcal{H}})$ the subset of operators with vanishing trace, ${\sf B}^{+}({\mathcal{H}})\subset{\sf B}({\mathcal{H}})$ the cone consisting of positively semi-definite operators, ${\sf B}^{+}_{1}({\mathcal{H}})\subset{\sf B}^{+}({\mathcal{H}})$ the convex set of statistical operators. and ${\sf U}({\mathcal{H}})$ the group of unitary operators acting on $\mathcal{H}$}. There exists a canonical isomorphism $\imath:{\mathcal{A}}\otimes{\mathcal{A}}\rightarrow L({\mathcal{H}}\otimes{\mathcal{H}})$ given by
@@
$\langle{\mathcal{X}}\left|\right.{\mathcal{Y}}\rangle:=\mbox{Tr}\left({\mathcal{X}}^{\ast}\,{\mathcal{Y}}\right),\;{\mathcal{X}},{\mathcal{Y}}\in L({\mathcal{A}})\;.$ (4)
@@
We will formulate the definitions concerning “complete positivity" only for the special case of a linear map $\phi:{\mathcal{A}}\rightarrow{\mathcal{A}}$ considered in this article, following NC00 and C06 . Recall that $L\left({\mathbb{C}}^{M}\right)$ denotes the algebra of complex $M\times M$-matrices. $\phi$ can be extended to $\phi\otimes{\mathbbm{1}}_{N}:{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)\rightarrow{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)$ by $(\phi\otimes{\mathbbm{1}})(A\otimes M):=\phi(A)\otimes M$ and linear extension. Then $\phi$ is called “completely positive" iff $X\geq 0$ implies $(\phi\otimes{\mathbbm{1}}_{N})(X)\geq 0$ for all Hermitean operators $X\in{\mathcal{A}}\otimes L\left({\mathbb{C}}^{M}\right)$. This condition is stronger than the property that $\phi$ maps positively semi-definite operators onto positively semi-definite ones. The usual example to illustrate this fact is $\tau(A)=A^{\top}$ (transposition). $\tau$ leaves the eigenvalues of any positively semi-definite operator invariant but is not completely positive: Set $N=2$ and ${\mathcal{A}}=L\left({\mathbb{C}}^{2}\right)$ and consider the matrix $X\in{\mathcal{A}}\otimes L\left({\mathbb{C}}^{2}\right)$ given by
@@
with eigenvalues $(2,0,0,0)$. $\tau\otimes{\mathbbm{1}}_{2}$ swaps the two anti-diagonal blocks of $X$
$(\tau\otimes{\mathbbm{1}}_{2})(X)=\left(\begin{array}[]{cc|cc}1&0&0&0\\ 0&0&1&0\\ \hline\cr 0&1&0&0\\ 0&0&0&1\\ \end{array}\right)\;,$ (6)
and thus yields a matrix with eigenvalues (1,1,1,1)(-1,1,1,1).
@@ IV.1 Choi isomorphism as a special case of the GKS isomorphism
It is shown in GKS76 that this family actually defines a basis of $L({\mathcal{A}})$ but we will not use this fact at the moment and postpone its proof. Let us first focus on the special case that an orthonormal basis $\left(|i\rangle\right)_{0\leq i<N}$ of $\mathcal{H}$} is chosen and consider the derived orthonormal basis $\left(|j\rangle\,\langle i|\right)_{0\leq i,j<N}$ of $\mathcal{A}$}. We write $\alpha\equiv(i,j)$ as a double index and set $E_{\alpha}:=|j\rangle\,\langle i|$, hence
$E_{\alpha}^{\ast}:=|i\rangle\,\langle j|\quad\mbox{for }0\leq i,j<N\;.$ (20)
Note the reversed positions of the indices. Then it can be shown that
@@
$\left\langle{\mathcal{E}}_{\alpha\beta}\left|\right.{\mathcal{E}}_{\gamma\delta}\right\rangle$ $=$ $\mbox{Tr }\left({\mathcal{E}}_{\alpha\beta}^{\ast}\,{\mathcal{E}}_{\gamma\delta}\right)$ (21)
@@
$=$ $\sum_{\epsilon}\mbox{Tr }\left(E_{\beta}\,E_{\epsilon}^{\ast}\,E_{\alpha}^{\ast}\,E_{\gamma}\,E_{\epsilon}\,E_{\delta}^{\ast}\right)$ (25)
@@
Then the matrix with entries $\widetilde{c}^{\alpha\beta}$ is exactly the Choi matrix of $\mathcal{E}$}. More precisely, we have
@@ IV.2 Definition of the GKS isomorphisms
    $=$ $\sum_{\alpha^{\prime}\,\beta^{\prime}}\overline{U_{\alpha\alpha^{\prime}}}\,U_{\beta\beta^{\prime}}\,\delta_{\alpha^{\prime}\,\beta^{\prime}}=\sum_{\alpha^{\prime}}\overline{U_{\alpha\alpha^{\prime}}}\,U_{\beta\alpha^{\prime}}=\sum_{\alpha^{\prime}}U_{\beta\alpha^{\prime}}\,U^{\ast}_{\alpha^{\prime}\,\alpha}\;.$ (38)
@@
$G$ $:$ $L({\mathcal{A}})\rightarrow L({\mathbb{C}}^{N\times N}),$ (45)
@@
The comparison of (43) and (29) shows that the Choi matrix isomorphism is a special case of the GKS-isomorphism. Already the Choi matrix isomorphism is not a single mathematical object but rather a family since it depends on a family of orthonormal bases in $\mathcal{A}$} that are of the form $\left(|i\rangle\langle j|\right)_{0\leq i,j<N}$ w. r. t. the family of orthonormal bases in $\mathcal{H}$} denoted by $\left(|i\rangle\right)_{0\leq i<N}$. The GKS-isomorphism $G$ further extends this family by considering general orthonormal bases $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in $\mathcal{A}$}. Of course, it still has to be shown that $G$ has analogous properties as the Choi matrix isomorphism.
@@ IV.3 Properties of the GKS isomorphisms
$=:$ $\sum_{\alpha^{\prime}\beta^{\prime}}\hat{g}^{\alpha^{\prime}\beta^{\prime}}\,G_{\alpha^{\prime}}\,A\,G_{\beta^{\prime}}^{\ast}\;.$ (49)
@@
[…]
@@
$=:$ $\sum_{\gamma}g^{\gamma}\,A_{\gamma}\,X\,A_{\gamma}^{\ast}\;.$ (58)
@@ V Comparison with other isomorphisms
There exist various other generalizations of the Choi isomorphism and the question arises whether the GKS-isomorphism considered above is really different. We will check this by an application to the example of the transposition $\tau$ of $2\times 2$-matrices considered in Section II which maps positive matrices to positive ones but is not completely positive.
@@
in $\mathcal{A}$}. Since $\tau$ only swaps $E_{1}$ and $E_{2}$ we can represent it by the matrix
$\widetilde{\tau}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{array}\right)\;.$ (61)
According to (17) the Choi matrix $c(\tau)$ of $\tau$} is obtained by a certain permutation of matrix elements, see Figure 1. It follows that $c(\tau)=\widetilde{\tau}$ and hence $c({\tau})$ is not positively semi-definite in accordance with the fact that $\tau$ is not completely positive.
@@
in $\mathcal{A}$}. After some calculations we obtain the following GKS-matrix $G(\tau)$
$G(\tau)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&1\\ \end{array}\right)\;,$ (63)
@@ V.1 Comparison with dePillis-Jamiołski isomorphism
As already noted in PS13 and again in FC24 , the isomorphism due to dP67 and J72 differs from the Choi isomorphism, although this difference is sometimes blurred by the misleading denotation “Choi-Jamiołski isomorphism". It can be written in the form
@@
seemingly depending on an orthonormal basis $\left(F_{\alpha}\right)_{0\leq\alpha<N^{2}}$ in $\mathcal{A}$}, but $J({\mathcal{E}})$ can be shown to be independent of this basis, see FC24 . Again we calculate the matrix of $J(\tau)$ where $\tau$ denotes the transposition of $2$-matrices and obtain
$J(\tau)=\left(\begin{array}[]{cccc}1&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ 1&0&0&1\\ \end{array}\right)\;,$ (65)